# ranDecepter: Real-time Identification and Deterrence of Ransomware Attacks

### v2

@@ III-B Realtime Phase: RW Identification using Hooks
Stage 2: If no suspicious file creation is detected, the system enters Stage 2. Here, the system evaluates whether the process possesses encryption capability. This is determined using three orthogonal techniques introduced in Section III-A: (1) detection of calls to standard cryptographic APIs like CryptEncrypt, AES_encrypt, or AES::Encryption; (2) runtime memory scanning for Cryptographic Function Signatures (CFS) when WriteFile is called; and (3) entropy-based heuristics using tools like ent and Volatility.
When standard crypto API invocations are intercepted, the system turns on the Encryption Flag. For embedded encryption routines, the CFS check is triggered during WriteFile, targeting the memory buffer just before data is flushed to disk—where encryption routines typically complete. This precision avoids full memory scans and ensures lightweight operation. If encryption capability is confirmed and ARC is enabled, the system returns FakeSuccess responses to RW for all sensitive operations from this point forward. If ARC is disabled, encryption is allowed to proceed, with additional safeguards activated in the following stages.
@@ III-C Reset Phase: Depleting Attackers’ Resources
To implement this looping mechanism, ranDecepter must first determine where the malware completes its “Transfer Key” routine—the stage where encryption keys and system-specific identifiers are sent to the attacker. This is accomplished through a customized dynamic analysis module, the Symbolic Binary Analyzer (SBA). Built on top of gExtractor [4], SBA uses selective symbolic execution to record user-space API invocations (over 390 supported API functions), log argument values, and capture the call stack state for each function, including the caller’s address and the API’s return target. It also identifies the binary’s entry point address, which is essential for loop redirection.
@@ IV Evaluations
# of RW Samples Identification Stage Files Lost 204 CreateFile with known RW extension 0 895 Encryption, File deletion or overwrite, Ransom note 0 16 High entropy Avg: 7 19 Never reached encryption (a) Detection Stages vs. File Loss    Application Overhead (%) 7zip 5.73 Firefox 5.44 WinSCP 4.61 Opera 3.79 MySQL Workbench 3.69 (b) Application Overhead Sample Extra DB Entries Extra Space (MB) (1h / 12h / 24h) (1h / 12h / 24h) r1 290K / 3420K / 6910K 43 / 527 / 1054 r2 131K / 1582K / 3081K 15 / 189 / 385 r3 413K / 4829K / 9223K 52.9 / 603 / 1189 r4 195K / 2257K / 4357K 22.7 / 272 / 544 (c) DB Misinformation Depletion
TABLE IV: Summary of (a) RW detection stages and file loss, (b) runtime overhead of common applications, and (c) attacker-side resource depletion due to deceptive DB entries.
@@ IV-A Evaluation of Accuracy and Effectiveness against RW
We evaluated ranDecepter’s detection accuracy and overall effectiveness using the 1,134 RW samples from our dataset. Of these, 204 samples were detected early during CreateFile operations by matching known RW file extensions, enabling immediate termination before any damage. The majority (895 samples) delayed renaming until after encryption, using MoveFile or MoveFileWithProgress. For these, ranDecepter leveraged MSG and CFS knowledge to flag suspicious encryption behavior, intercepting and faking subsequent DeleteFile and WriteFile operations via the FakeSuccess strategy to preserve original files. Ransom note creation confirmed RW behavior, triggering alerts with zero data loss. Among the remaining 35 samples, 19 opened network ports but remained inactive due to missing C&C triggers, while 16 employed custom, embedded crypto routines. These were eventually flagged by entropy-based heuristics, though with an average loss of seven files. Overall, ranDecepter achieved 100% identification accuracy, effectively detecting RW using static, dynamic, and custom cryptographic techniques. Results are summarized in Table IV(a).
@@ IV-B Evaluation against Benign Applications
During evaluation, we executed the benign applications and monitored for false positives, with particular attention to behaviors involving encryption and file deletion or renaming. As expected, no application was misclassified as RW during the CreateFile stage. While several applications performed encryption or file operations, they were not flagged as malicious, since they did not create files with known RW extensions or generate ransom notes. To assess performance impact, we measured each application’s response time with and without our system’s API hooking. All measurements were conducted using WinAppDriver to ensure consistency and automation. As summarized in Table IV(b), the maximum observed overhead was 5.73% for 7Zip, followed by 5.44% for Mozilla Firefox, 4.61% for WinSCP, 3.79% for Opera, and 3.69% for MySQL Workbench. Due to space constraints, the top five applications with the highest overheads are shown; the complete evaluation results for all 45 applications are available in [5]. The remaining applications experienced overheads ranging from 0.40% to 2.99%, demonstrating that ranDecepter imposes minimal performance degradation in practice.
@@ IV-C Comparison with Existing Similar Work
TABLE V: Comparison of ranDecepter with Existing RW Defenses
RWGuard uses decoy file monitoring, CryptoAPI hooking, and a process-level anomaly detector based on I/O activity. However, its reliance on standard APIs and coarse behavioral metrics results in significant file loss, averaging 288 legitimate files encrypted before detection. In contrast, ranDecepter prevents file loss in 98.5% of cases and limits it to just 7 files in the remaining 1.5%, without requiring model training or process profiling. R-Sentry and R-Trap both employ decoy files. R-Sentry places honeyfiles based on assumed directory traversal, while R-Trap uses machine learning to adaptively clone user files. Although more resilient than static decoys, both suffer false positives when benign applications access decoys. R-Sentry allows up to 10 files to be lost before alerting, and R-Trap averages 18. R-Trap also incurs moderate overhead from continuous decoy regeneration and ML-driven selection. In contrast, ranDecepter requires no decoy or ML infrastructure, avoids false positives, and achieves near-zero file loss with lightweight operation. R-Locker introduces a FIFO-based trap file in user directories, terminating any process that accesses it. This assumes RW follows breadth-first traversal and encounters the FIFO early, which may not hold for targeted RW. Like other decoy-triggered systems, it lacks confirmation logic, increasing the risk of benign interference. Table V summarizes these systems across key deployment metrics: false positives, operational overhead, and file loss.
@@ IV-D Accuracy and Performance Analysis of Binary Orchestration in the Reset Phase
Each sample was compiled and analyzed within our Symbolic Binary Analyzer (SBA) to extract encryption-relevant API addresses. The Actuating Agent then generated the corresponding Deceptor DLL configurations and injected them into the malware processes. As the binaries executed, server logs confirmed repeated receipt of unique victim IDs and encryption keys, confirming 100% accuracy in binary orchestration. For performance assessment, we monitored the attacker’s database over 24 hours to assess resource depletion. Our system generated 3,081K to 9,223K entries, consuming 385 MB to 1.1 GB of memory. Using VPNs/proxies could further enhance this by masking traffic origins. Results are summarized in Table IV(c).
@@ V Related work
Deception-Based Methods: Deception techniques attempt early detection by populating the filesystem with decoy to lure RW activity [21, 28]. These approaches, while conceptually proactive, face several challenges. First, they require wide deployment of synthetic artifacts, which may overlap with legitimate user workflows, introducing false positives [15, 29]. Second, adaptive RW can selectively avoid decoys, targeting specific files. To address these limitations, recent systems such as R-Trap [14] incorporate metadata-aware cloning and ML-driven decoy placement to enhance realism. However, these remain reactive, as action is only triggered after partial encryption has occurred. For instance, RWGuard, R-Sentry, and R-Trap collectively permit between 10 and 288 file losses before mitigation [21, 14, 28]. Furthermore, all these techniques depend on file system-level indicators, limiting scalability and exposing them to evasion via metadata manipulation.
@@ VI Discussion and Conclusion
While our approach demonstrates promising outcomes, certain limitations must be acknowledged. Advanced RW may bypass user-space API hooks by verifying or unhooking IAT/EAT entries. To counter this, we plan to incorporate more resilient techniques such as runtime obfuscation, kernel-level hooks (e.g., SSDT, callbacks) [17], and execute-only memory protections using Extended Page Tables (EPT). These methods make hook detection more difficult but require careful tuning to avoid performance overhead. Although our resource depletion strategy raises the attacker’s operational costs, sophisticated adversaries may adapt. Nonetheless, the goal extends beyond storage exhaustion by enabling proactive measures like beaconing to reveal C&C infrastructure, automated sinkholing, and malware poisoning. These tactics provide both disruption and actionable intelligence for law enforcement. We also acknowledge possible unintended consequences, such as victims losing access if attacker infrastructure is dismantled post-payment. However, our focus is on disrupting active RW campaigns, consistent with botnet takedown strategies. Additionally, ranDecepter currently emphasizes parent-child process monitoring. This intra-process focus may overlook RW coordinating across independent processes. Future work will explore inter-process correlation techniques, such as policy-based process correlation [8], and lineage analysis, to enhance detection of distributed RW behavior.

### v3

@@ III-B Realtime Phase: RW Identification using Hooks
Stage 2: If no suspicious file creation is detected, the system enters Stage 2. Here, the system evaluates whether the process possesses encryption capability. This is determined using three orthogonal techniques introduced in Section III-A: (1) detection of calls to standard cryptographic APIs like CryptEncrypt, AES_encrypt, or AES::Encryption; (2) runtime memory scanning for Cryptographic Function Signatures (CFS) when WriteFile is called; and (3) entropy-based heuristics using tools like ent and Volatility. When standard crypto API invocations are intercepted, the system turns on the Encryption Flag. For embedded encryption routines, the CFS check is triggered during WriteFile, targeting the memory buffer just before data is flushed to disk—where encryption routines typically complete. This precision avoids full memory scans and ensures lightweight operation. If encryption capability is confirmed and ARC is enabled, the system returns FakeSuccess responses to RW for all sensitive operations from this point forward. If ARC is disabled, encryption is allowed to proceed, with additional safeguards activated in the following stages.
@@ III-C Reset Phase: Depleting Attackers’ Resources
To implement the looping mechanism, ranDecepter must first locate the malware’s “Transfer Key” stage, where encryption keys and system identifiers are sent to the attacker. This is achieved using the Symbolic Binary Analyzer (SBA), a custom dynamic analysis module built on gExtractor [6, 29]. SBA employs selective symbolic execution to trace over 390 user-space API calls, log argument values, and capture call stack state, including the caller’s address and API return target. It also extracts the binary’s entry point, critical for redirecting execution into the loop.
@@ IV-A Evaluation of Accuracy and Effectiveness against RW
We evaluated ranDecepter’s detection accuracy and overall effectiveness using the 1,134 RW samples from our dataset. Of these, 204 samples were detected early during CreateFile operations by matching known RW file extensions, enabling immediate termination before any damage. The majority (895 samples) delayed renaming until after encryption, using MoveFile or MoveFileWithProgress. For these, ranDecepter leveraged MSG and CFS knowledge to flag suspicious encryption behavior, intercepting and faking subsequent DeleteFile and WriteFile operations via the FakeSuccess strategy to preserve original files. Ransom note creation confirmed RW behavior, triggering alerts with zero data loss. Among the remaining 35 samples, 19 opened network ports but remained inactive due to missing C&C triggers, while 16 employed custom, embedded crypto routines. These were eventually flagged by entropy-based heuristics, though with an average loss of seven files. Overall, ranDecepter achieved 100% identification accuracy, effectively detecting RW using static, dynamic, and custom cryptographic techniques. Results are summarized in Table IV.
# of RW Samples Identification Stage Files Lost
204 CreateFile with known RW extension 0
895 Encryption, File deletion or overwrite, Ransom note 0
16 High entropy Avg: 7
19 Never reached encryption
TABLE IV: Detection Stages vs. File Loss
@@ IV-B Evaluation against Benign Applications
During evaluation, we executed the benign applications and monitored for false positives, with particular attention to behaviors involving encryption and file deletion or renaming. As expected, no application was misclassified as RW during the CreateFile stage. While several applications performed encryption or file operations, they were not flagged as malicious, since they did not create files with known RW extensions or generate ransom notes. To assess performance impact, we measured each application’s response time with and without our system’s API hooking. All measurements were conducted using WinAppDriver to ensure consistency and automation. As summarized in Table V, the maximum observed overhead was 5.73% for 7Zip, followed by 5.44% for Mozilla Firefox, 4.61% for WinSCP, 3.79% for Opera, and 3.69% for MySQL Workbench. Due to space constraints, the top ten applications with the highest overheads are shown; the complete evaluation results for all 45 applications are available in [1]. The remaining applications experienced overheads ranging from 0.40% to 2.99%, demonstrating that ranDecepter imposes minimal performance degradation in practice.
Application Overhead (%) Application Overhead (%)
7zip 5.73 WinRAR 3.41
Firefox 5.44 Picasa 3.19
WinSCP 4.61 Calibre 3.14
Opera 3.79 Foxit PDF Reader 3.08
MySQL Workbench 3.69 Photoshop 3.06
TABLE V: Runtime Overhead of Common Applications
@@ IV-C Comparison with Existing Similar Work
TABLE VI: Comparison of ranDecepter with Existing RW Defenses
RWGuard uses decoy file monitoring, CryptoAPI hooking, and a process-level anomaly detector based on I/O activity. However, its reliance on standard APIs and coarse behavioral metrics results in significant file loss, averaging 288 legitimate files encrypted before detection. In contrast, ranDecepter prevents file loss in 98.5% of cases and limits it to just 7 files in the remaining 1.5%, without requiring model training or process profiling. R-Sentry and R-Trap both employ decoy files. R-Sentry places honeyfiles based on assumed directory traversal, while R-Trap uses machine learning to adaptively clone user files. Although more resilient than static decoys, both suffer false positives when benign applications access decoys. R-Sentry allows up to 10 files to be lost before alerting, and R-Trap averages 18. R-Trap also incurs moderate overhead from continuous decoy regeneration and ML-driven selection. In contrast, ranDecepter requires no decoy or ML infrastructure, avoids false positives, and achieves near-zero file loss with lightweight operation. R-Locker introduces a FIFO-based trap file in user directories, terminating any process that accesses it. This assumes RW follows breadth-first traversal and encounters the FIFO early, which may not hold for targeted RW. Like other decoy-triggered systems, it lacks confirmation logic, increasing the risk of benign interference. Table VI summarizes these systems across key deployment metrics: false positives, operational overhead, and file loss.
@@ IV-D Accuracy and Performance Analysis of Binary Orchestration in the Reset Phase
Each sample was compiled and analyzed within our Symbolic Binary Analyzer (SBA) to extract encryption-relevant API addresses. The Actuating Agent then generated the corresponding Deceptor DLL configurations and injected them into the malware processes. As the binaries executed, server logs confirmed repeated receipt of unique victim IDs and encryption keys, confirming 100% accuracy in binary orchestration. For performance assessment, we monitored the attacker’s database over 24 hours to assess resource depletion. Our system generated 3,081K to 9,223K entries, consuming 385 MB to 1.1 GB of memory. Using VPNs/proxies could further enhance this by masking traffic origins. Results are summarized in Table VII.
Sample Extra DB Entries Extra Space (MB)
(1h / 12h / 24h) (1h / 12h / 24h)
r1 290K / 3420K / 6910K 43 / 527 / 1054
r2 131K / 1582K / 3081K 15 / 189 / 385
r3 413K / 4829K / 9223K 52.9 / 603 / 1189
r4 195K / 2257K / 4357K 22.7 / 272 / 544
TABLE VII: Attacker-Side DB Misinformation Depletion
@@ V Related work
Deception-Based Methods: Deception techniques attempt early detection by populating the filesystem with decoy to lure RW activity [22, 31]. These approaches, while conceptually proactive, face several challenges. First, they require wide deployment of synthetic artifacts, which may overlap with legitimate user workflows, introducing false positives [16, 32, 3]. Second, adaptive RW can selectively avoid decoys, targeting specific files. To address these limitations, recent systems such as R-Trap [15] incorporate metadata-aware cloning and ML-driven decoy placement to enhance realism. However, these remain reactive, as action is only triggered after partial encryption has occurred. For instance, RWGuard, R-Sentry, and R-Trap collectively permit between 10 and 288 file losses before mitigation [22, 15, 31]. Furthermore, all these techniques depend on file system-level indicators, limiting scalability and exposing them to evasion via metadata manipulation.
@@ VI Discussion and Conclusion
While our approach demonstrates promising outcomes, certain limitations must be acknowledged. Advanced RW may bypass user-space API hooks by verifying or unhooking IAT/EAT entries. To counter this, we plan to incorporate more resilient techniques such as runtime obfuscation [26], kernel-level hooks [18], and execute-only memory protections using Extended Page Tables (EPT). These methods make hook detection more difficult but require careful tuning to avoid performance overhead. Although our resource depletion strategy raises the attacker’s operational costs, sophisticated adversaries may adapt. Nonetheless, the goal extends beyond storage exhaustion by enabling proactive measures like beaconing to reveal C&C infrastructure, automated sinkholing, and malware poisoning. These tactics provide both disruption and actionable intelligence for law enforcement. We also acknowledge possible unintended consequences, such as victims losing access if attacker infrastructure is dismantled post-payment. However, our focus is on disrupting active RW campaigns, consistent with botnet takedown strategies. Additionally, ranDecepter currently emphasizes parent-child process monitoring. This intra-process focus may overlook RW coordinating across independent processes. Future work will explore inter-process correlation techniques, such as policy-based process correlation [9], and lineage analysis, to enhance detection of distributed RW behavior.